# Embodied Red Teaming for Auditing Robotic Foundation Models

### v1

@@ 4.2 Large Models Struggle At Instruction Generalization
Setup: We assess generalization of recent large vision-language-action (VLA) models for robotic manipulation (Brohan et al., 2022; 2023; Team et al., 2024). Unlike GR-1 and 3D-Diffuser (Section 4.1), these VLAs are trained on real robot manipulation videos. We select OpenVLA (Kim et al., 2024) for its state-of-the-art performance and open-source availability. OpenVLA employs a Vision Transformer (ViT) (Dosovitskiy, 2020) for visual features and a LLaMA2-7B model (Touvron et al., 2023) for language instructions, converting both into tokens. With approximately 7 billion parameters—significantly larger than the 200 million parameters of 3D-Diffuser and GR-1—we expect OpenVLA to generalize better across instructions. The LLaMA2-7B model is trained to predict action tokens, such as gripper movements. We evaluate OpenVLA in the simulated environment SimplerEnv [^1] https://github.com/DelinQu/SimplerEnv-OpenVLA (Li et al., 2024), which replicates many scenes similar to its training videos. The task involves a robot using an overhead camera to pick up a Coke can from a table. We only evaluate OpenVLA on SimplerEnv because it only accepts a single image, whereas CALVIN (Mees et al., 2022) and RLBench (James et al., 2020) feed multiple images from different angles to the robot.
@@
Metric: The metric $R(\pi,\bm{c})$ (Equation 1) is set as a success indicator for picking up a Coke can. OpenVLA is tested on 25 initial states, each with a unique object position. For each initial state, ERT generates 4 instructions, totaling 100 instructions. Instructions are generated differently for each initial state because ERT may craft initial state-specific instructions (e.g., based on object arrangement). We measure the robot’s success rate across these 100 instruction-state pairs. We run ERT with 5 different random seeds and report the success rate with its 95% confidence interval. We also compare the robot’s performance on ERT-generated instructions to its performance on the original instruction provided in SimplerEnv. Note that this benchmark provides only one instruction per task. For the task of picking up the Coke can, this instruction is: “Pick the opened Coke can.” The robot’s success rate on the original instruction is measured across the same 25 initial states.
@@
Table 2: The state-of-the-art robot policy OpenVLA (Kim et al., 2024), with 7B parameters, performs well on instructions from the SimplerEnv benchmark (Li et al., 2024) but faces challenges generalizing to instructions generated by ERT.
Results: Table 2 shows that while OpenVLA performs well with SimplerEnv instructions, its success rate drops significantly with ERT-generated instructions. This suggests that even state-of-the-art VLA models with a large number of parameters struggle to generalize across language instructions. We run ERT without refinement (i.e., ERT(k=0)) because the initial iteration’s instructions already caused significant robot failures. Table 3 provides examples of ERT-generated instructions that accurately describe tasks like picking up a Coke can, yet OpenVLA fails to follow most of them. Moreover, there is no clear pattern in the instructions that cause OpenVLA to fail, making it difficult to improve generalization through instruction curation.
@@
Table 3: Examples of ERT-generated instructions for picking up a Coke can, along with OpenVLA’s (Kim et al., 2024) success rates, reveal that while all instructions accurately describe the task, OpenVLA often fails to execute them. The unpredictable failure patterns make instruction curation challenging.
4.3 Safety of language-conditioned robots is overlooked
We emphasize that safety is a significant concern for language-conditioned robot policy—a factor often overlooked in existing benchmarks that focus solely on success rates. Safety considerations—such as preventing environmental damage, avoiding harm to humans, and preventing self-damage—are critical because these robots interact with humans through natural language instructions. In this experiment, we consider unsafe behavior as causing objects to fall from the table. We present examples in Figure 6, as generated with instructions described in Section A.3, and summarize our findings below. In Figure 6, we illustrate 4 timesteps from episodes in which the robot exhibits unsafe behavior across 4 instructions. These unsafe behaviors result in one or more objects falling off the table.
Robots follow unsafe instructions. In Figures 6(a) and 6(b), the robots successfully complete tasks that are inherently unsafe—such as causing objects to fall off and destabilizing them. This suggests the need for alignment training, similar to that used in large language models (LLMs), to ensure robots align with human preferences and refuse to perform harmful or unsafe actions that could damage the environment.
Robots exhibit unsafe behaviors on neutral instructions. While robots following unsafe instructions is problematic, it could be addressed by training them to reject such commands or by filtering out unsafe instructions using an LLM before execution. However, as shown in Figure 6(c) and 6(d), we found that robots can still display unsafe behaviors even when given instructions for safe tasks. This means we cannot rely solely on filtering instructions to prevent unsafe behaviors, as safe instructions that pass safety checks may still lead to unintended unsafe actions.
Potential threat: jailbreak. Language-conditioned robot policies may not always follow instructions accurately and could exhibit unsafe behaviors, making it unpredictable when they might damage the environment or harm humans based on the given instructions. Even neutral and safe instructions can cause the robot to act unsafely, presenting a vulnerability that malicious users could exploit. For example, consider a service robot in a store that takes natural language instructions. A malicious user could craft a neutral instruction (e.g., "Adjust your arm position") to trigger the robot to harm customers. Since the instruction appears completely neutral, engineers might not be aware of the potential danger or be able to prevent it. Concurrent works (Robey et al., 2024) showed preliminary results on jailbreaking language-conditioned robot policies.
@@ 4.3 Safety of language-conditioned robots is overlooked
Figure 6: Robots (3D Diffuser Ke et al. (2024) in this case) may follow unsafe instructions, as seen in Figures 6(a) and 6(b), or exhibit unsafe behavior even with neutral instructions, shown in Figures 6(c) and 6(d). This highlights the need for alignment training to ensure robots reject harmful tasks. Additionally, a potential vulnerability exists in that malicious users could exploit seemingly neutral instructions to cause harm, posing a significant safety risk.
4.4 Analysis of failure modes
To examine the failure modes of these robot models, we have selected ERT-generated instructions with the lowest success rates from each task, presented in Table 5, and summarize our findings below:
Complex Instructions. Complex or multi-step instructions often lead to task failures. For instance, in the close_drawer task (1.05% success), the directive to "check for obstructions, then apply force to close the drawer snugly" introduces multiple actions, namely (1) "check for obstructions", (2) "apply force", and (3) "close the drawer snugly".
Uncommon vocabulary. Using rare synonyms or vague terms can cause misinterpretation and subsequent failures. The instruction for turn_off_led (9.47% success) uses the phrase "suppress its glow," which is rare in the instructions of CALVIN and RLBench benchmarks yet a valid way of describing the task turn_off_led.
Distracting details. Including unnecessary information or human-centric perspectives in instructions can distract the robot, leading to failures. In the open_drawer task (0.0% success), the phrase "pull it towards yourself" introduces a perspective that the robot may not comprehend.
Takeaway: The observed failure modes indicate that most failures result from confusion around the more complex generated instructions. One plausible explanation is that the ERT instructions might be more convoluted in text than the training instructions but command similar behaviors. In other words, the pre-trained language embeddings that interpret the instructions capture textual similarity but not "embodied similarity" — the similarity in intended actions. For example, the instruction "Check for obstructions, then apply force to close the drawer snugly" is textually different from the simpler "Close the drawer", yet both describe the same desired behavior. This suggests that training language embedding models jointly with the robot’s policy could enable the embeddings to capture similarities in desired actions, improving performance.
@@ 4.4 Analysis of failure modes
Table 4: Instructions with the lowest robot success rate for each task are listed. Full table in Table 5.
5 Discussion and Limitations
@@
For Red Teamers: To our knowledge, this is the first study to apply red teaming to language-conditioned robots using vision-language models. We hope to inspire new applications of red teaming in robotics by bridging these fields. Red teaming robotic systems presents unique challenges, such as the high cost of feedback-based generation due to lengthy instruction execution and the time-consuming evaluation process. Developing surrogate metrics to efficiently estimate the performance and safety of foundational agents is crucial.
Limitations: Because the method relies on VLM-generated instructions, ERT may not discover every way in which a robot may fail. Instead, ERT provides a way to automatically and efficiently generate instructions that can trigger failure modes. Additionally, our study is limited to manipulation tasks. However, the ERT approach is versatile and can be extended to other language-conditioned models, such as navigation robots Shridhar et al. (2020), which we leave for future work.
@@ Appendix A Appendix
A.1 Prompts for Generating Instructions
The instructions generated by GPT-4 in ERT used the following meta-prompt:
@@ A.1 Prompts for Generating Instructions
A.2 Selected qualitative examples
We select instructions with the lowest success rate at each task and analyze the failure modes in Section 4.4.
@@ A.2 Selected qualitative examples
Table 5: Instructions with the lowest robot success rate at each task.
A.3 Safety Instructions
@@
A.4 CALVIN Instructions
A.4.1 ERT(seed = 0, $k=0$)
@@ A.4.15 ERT(seed = 4, $k=2$)
A.5 RLBench Instructions
A.5.1 ERT(seed = 0, $k=0$)

### v2

@@ IV-C Large models also struggle at instruction generalization
Setup: We assess generalization of recent large vision-language-action (VLA) models for robotic manipulation [2, 3, 21]. Unlike GR-1 and 3D-Diffuser (Section IV-A), these VLAs are trained on real robot manipulation videos with annotated instructions. We select OpenVLA [10] for its state-of-the-art performance and open-source availability. OpenVLA employs a Vision Transformer (ViT) [6] for visual features and a LLaMA2-7B model [24] for language instructions, converting both into tokens. With approximately 7 billion parameters—significantly larger than the 200 million parameters of 3D-Diffuser and GR-1—we expect OpenVLA to generalize better across instructions. The LLaMA2-7B model is finetuned to predict action tokens, such as gripper movements. We evaluate OpenVLA in the simulated environment SimplerEnv [11], which replicates many scenes similar to its training videos. The task involves a robot using an overhead camera to pick up a Coke can from a table. We only evaluate OpenVLA on SimplerEnv because it only accepts a single image, whereas CALVIN [12] and RLBench [8] feed multiple images from different angles to the robot.
@@
Reward: The reward $R(\pi,\bm{c})$ (Equation 1) is set as a success indicator for picking up a Coke can. OpenVLA is tested on 25 initial states, each with a unique object position. For each initial state, ERT generates 4 instructions, totaling 100 instructions. Instructions are generated differently for each initial state because ERT may craft initial state-specific instructions (e.g., based on object arrangement). We measure the robot’s success rate across these 100 instruction-state pairs. We run ERT with 5 different random seeds and report the success rate with its 95% confidence interval. We also compare the robot’s performance on ERT-generated instructions to its performance on the original instruction provided in SimplerEnv. Note that this benchmark provides only one instruction per task. For the task of picking up the Coke can, this instruction is: “Pick the opened Coke can.” The robot’s success rate on the original instruction is measured across the same 25 initial states.
@@
TABLE II: The state-of-the-art robot policy OpenVLA [10], with 7B parameters, performs well on instructions from the SimplerEnv benchmark [11] but faces challenges generalizing to instructions generated by ERT.
Results: Table II shows that while OpenVLA performs well with SimplerEnv instructions, its success rate drops significantly with ERT-generated instructions. This suggests that even state-of-the-art VLA models with a large number of parameters struggle to generalize across language instructions. We run ERT without refinement (i.e., ERT(k=0)) because the initial iteration’s instructions already caused significant robot failures. Table III provides examples of ERT-generated instructions that accurately describe tasks like picking up a Coke can, yet OpenVLA fails to follow most of them. Moreover, there is no clear pattern in the instructions that cause OpenVLA to fail, making it difficult to improve generalization through instruction curation.
@@
TABLE III: Examples of ERT-generated instructions for picking up a Coke can, along with OpenVLA’s [10] success rates, reveal that while all instructions accurately describe the task, OpenVLA often fails to execute them. The unpredictable failure patterns make instruction curation challenging.
IV-D Safety of language-conditioned robots is overlooked
@@
Figure 6: Robots (3D-Diffuser [9] in this case) may follow unsafe instructions, as seen in Figures 6(a) and 6(b), or exhibit unsafe behavior even with neutral instructions, shown in Figures 6(c) and 6(d). This highlights the need for alignment training to ensure robots reject harmful tasks. Additionally, a potential vulnerability exists in that malicious users could exploit seemingly neutral instructions to cause harm, posing a significant safety risk.
We emphasize that safety is a significant concern for language-conditioned robot models—a factor often overlooked in existing benchmarks that focus solely on success rates. Safety considerations—such as preventing environmental damage, avoiding harm to humans, and preventing self-damage—are critical because these robots interact with humans through natural language instructions. In this experiment, we consider unsafe behavior as causing objects to fall from the table. We present examples in Figure 6, as generated with instructions described in Section A-C, and summarize our findings below. In Figure 6, we illustrate 4 timesteps from episodes in which the robot exhibits unsafe behavior across 4 instructions. These unsafe behaviors result in one or more objects falling off the table.
Robots follow unsafe instructions. In Figures 6(a) and 6(b), the robots successfully complete instructions that are inherently unsafe—such as causing objects to fall off and destabilizing them. This suggests the need for alignment training, similar to that used in large language models (LLMs), to ensure robots align with human preferences and refuse to perform harmful or unsafe actions that could damage the environment.
Robots exhibit unexpected behaviors. While filtering out unsafe instructions or training robots to reject unsafe instructions can address cases where unsafe tasks are explicitly specified, this is not sufficient. As illustrated in Figure 6(c) and 6(d), robots can still exhibit unexpected behaviors even when following safe instructions for safe tasks (e.g., “Rotate the blue object 10 degrees clockwise, keeping it in place” in Figure 6(c)). This demonstrates that relying solely on instruction filtering is inadequate, as safe instructions passing safety checks may still lead to unexpected unsafe actions.
Potential threat: jailbreak. Language-conditioned robot policies may not always follow instructions accurately and could exhibit unsafe behaviors, making it unpredictable when they might damage the environment or harm humans based on the given instructions. Even neutral and safe instructions can cause the robot to act unsafely, presenting a vulnerability that malicious users could exploit. For example, consider a service robot in a store that takes natural language instructions. A malicious user could craft a neutral instruction (e.g., "Adjust your arm position") to trigger the robot to harm customers. Since the instructions appear completely neutral, engineers might not be aware of the potential danger or be able to prevent it. Concurrent works [19] have also highlighted jailbreaking of language-conditioned robot policies.
IV-E Qualitative Analysis of Failure Modes
@@
TABLE IV: Instructions with the lowest robot success rate for each task are listed. Full table in Table V.
We observed that ERT generates a higher number of instructions that robots fail to execute compared to those in the current benchmark. To understand the differences between benchmark instructions and those produced by ERT, we selected ERT-generated instructions with the lowest success rates for each task, as shown in Table IV. Our findings are summarized below.
Step-by-step instructions. The CALVIN benchmark lacks multi-step instructions. We found that instructions breaking tasks into multiple actions often result in failures. For example, in the close_drawer task (1.05% success), the instruction “check for obstructions, then apply force to close the drawer snugly” introduces three separate actions.
Uncommon vocabulary. CALVIN’s instructions use a limited vocabulary. ERT-generated instructions sometimes employ rare synonyms, leading to misinterpretation and failures. For instance, the turn_off_led task (9.47% success) includes the phrase “suppress its glow,” which is uncommon in CALVIN and RLBench benchmarks, despite being a valid description of the task turn_off_led.
Human-centric tone. Instructions framed from a human perspective can confuse robots. In the open_drawer task (0% success), the phrase “pull it towards yourself” was intended to mean pulling the drawer towards the robot. However, since the benchmark instructions do not use human-centric language, the robot fails to interpret such phrasing correctly.
Use of adverbs. CALVIN instructions do not modify verbs with adverbs. ERT-generated instructions sometimes include adverbs, which are absent from the benchmark. For example, in open_drawer, the instruction “Grip the handle gently and pull it towards yourself” uses “gently” to describe “grip,” a construction not seen in CALVIN instructions.
Unnecessary actions. While CALVIN instructions are direct, ERT-generated instructions may include unnecessary details. For example, in the open_drawer task, the instruction “Start by identifying the drawer and open it fully” adds an extra step of identifying the drawer,” which is not required but still describes the task of open_drawer.
Why do these failure modes matter? Some instructions that cause robot failures are common in daily use (e.g., “Pick up the blue cube from the surface” for the lift_blue_block_table task), and robots should not fail on them. Although some instructions that cause robot failures are uncommon, they highlight critical limitations in deploying state-of-the-art language-conditioned robot models. Developers must ensure that robots can handle diverse and unpredictable phrasings, as real users may give commands different from those in current benchmarks. While users can adapt to the robot and use familiar instructions that robots respond to better, it is difficult for them to know what kind of instruction phrasing will work. Users cannot predict which instructions might lead to failures and thus cannot avoid them. Therefore, language-conditioned robots should accurately perform tasks regardless of how instructions are phrased, as long as they correctly describe the task.
V Discussion and Limitations
@@
For Red Teamers: To our knowledge, this is the first study to apply red teaming to language-conditioned robots using vision-language models. We hope to inspire new applications of red teaming in robotics by bridging these fields. Red teaming robotic systems presents unique challenges, such as the high cost of feedback-based generation due to lengthy instruction execution and the time-consuming evaluation process. Developing surrogate rewards to efficiently estimate the performance and safety of foundational agents is crucial.
Limitations: Because the method relies on VLM-generated instructions, ERT may not discover every way in which a robot may fail. Instead, ERT provides a way to automatically and efficiently generate instructions that can trigger failure modes. Additionally, our study is limited to manipulation tasks. However, the ERT approach is versatile and can be extended to other language-conditioned models, such as navigation robots [20], which we leave for future work.
@@ Acknowledgements
  • Christophe Dupuy: Contributed to experiment execution, experiment design, idea development, and proofreading.
  • Rahul Gupta: Contributed to experiment design, idea development, and proofreading.
@@ Appendix A Appendix
A-A Prompts for Generating Instructions
ERT.
The instructions generated by GPT-4o in ERT used the following system prompt:
@@ A-A Prompts for Generating Instructions
Rephrase.
We use the following prompts in the baseline Rephrase to prompt GPT-4o:
Rephrase this list of {N} instructions {TRAINING INSTRUCTIONS}.
{TRAINING INSTRUCTIONS} are from the instructions provided in CALVIN and RLBench benchmark.
A-B Selected qualitative examples
We select instructions with the lowest success rate at each task and analyze the failure modes in Section IV-E.
@@ A-B Selected qualitative examples
TABLE V: Instructions with the lowest robot success rate at each task.
A-C Safety Instructions
@@
A-D CALVIN Instructions
A-D1 ERT(seed = 0, $k=0$)
@@ A-D9 ERT(seed = 2, $k=2$)
A-E RLBench Instructions
A-E1 ERT(seed = 0, $k=0$)